You are here: Foswiki>Commissioning Web>Com20101220 (22 Dec 2010, RichardGreen)Edit Attach

RFBG Commissioning: UTC 2010 December 20

Observer(s): D. Miller
Telescope Operator: D. Gonzalez (LBTO)
Telescope Support: M. Midkiff (LBTO)
Instrument Support: J. Morris (LBTO)

Summary

Lost the entire night to bad weather (humidity, clouds, high winds). Since this is our 5th and final night of Run 3, we officially have a bona-fide shut-out, at least insofar as on-sky work is concerned.

Instead, we pushed through a lot of our nominal daytime list of tasks for instrument characterization, calibration and scripting procedures, and general exercising of the system and getting bugs out. We also implemented, tested, and refined the startup procedures that will be used by staff and [skilled] observers during normal operations, getting somewhat ahead on those tasks.

The instrument performed flawlessly with no glitches. All computer worked without problems, and were able to be restarted after the UPS failure, with one exception that is noted and will be tested back home (it was minor, but hard to diagnose - the mitigation was clear once recognized, but implementing that in a more convenient tool is what we need to test in the lab back at OSU).

Preparations

Dewars filled at 1615MST - still holding well.

A few software tests during the afternoon, but otherwise we're ready to go. With bad weather, we'll run tests like mask repeatability, camera focus sequences, etc.

Details

All times are UTC unless otherwise noted.

At 1330MST, still high humidity, but looks partly clear behind. Breezy with gusts up to 23m/s. Temperature still hovering around 0°C like most of last night.

2130 [Dec 19] - Slit mask repeatability test.

Telescope at Zenith, using the red science camera 1Kx1K ROI and the pinhole sieve mask. 20 repeats of mask in and out, with snaps in between. Run using maskrep.pro Also took 10 baseline images with the mask unmoving to measure the ground RMS just for the mask measurement.

2200 [Dec 19] - Slit Mask images, blue and red 3Kx3K ROI. Images with the flat lamp of all 3 segmented long-slit masks

  • LS5x60x0.6
  • LS5x60x1.2
  • LS5x60x0.3

Run using maskim.pro

0000 - Humidity 100%, Temp 0C, wind 15-20m/s out of the SSW. Currently inside a blowing cloud.

Repeating maskrep.pro with the telescope at Horizon after the LBC-B fill

0100 - High humidty and winds 15-20m/s out of WSW, and inside a blowing cloud (getting the flashlight light-saber effect big time).

0101 - Repeating maskrep.pro with the telescope at 45° elevation, rotAngle=0°

0123 - Repeating maskrep.pro at El=45° and rotAngle=+45°

0142 - Repeating maskrep.pro at El=45° and rotAnge=-45°

0300 - Humidity 100%, Temp 0°C, still sitting inside a clouds.

0400 - Wet and drippy with Humidity 100%, temp ~0°C, and wind 15m/s out of the SSW.

0500 - Wind 15-20m/s out of the SW, humidity still 100%, but temperatur has dropped to -1.5°C. The past hour we've had episodes of thinning of the cloud sitting over Mt. Graham, but wind has also gusted regularly into the warning zone. Not looking good for tonight.

0700 - Humidity still 100%, wind 15-20m/s out of the SW, with over-limit gusts. Temperature has dropped to -1.5°C the past 2 hours, we can see ice forming on the VATT all-sky camera. We are no longer in cloud (probably why the temp is dropping), and you can see stars and the moon and high clouds around us (consistent with the current IR satellite images). Ice is forming on the post for the solar tracker. Very windy on the patio.

0800 - Humidity has dropped to 98%, Temp is -2°C, and it was clear for a bit with the fog gone, but now the next waves of high cirrus are moving in. Wind has been gusting above 20m/s, so we'd be closed for wind if it was drier.

0900 - Humidity now 94%, Temp -2°C, but the wind is strong and gusty - the strongest was 26m/s in the past 10 minutes. Another band of clouds is encroaching from the NW and visible in the bright moonlight. Judging from the IR satellite images, they'll be over us by about 10UT.

1000 - We stuck it out until 3am, but the humidty and wind stayed above limits, and the telescope mirrors would be 6.5°C warmer than the air, so we'd have no chance of equilibrating enough to matter. Another wave of thick high cloud is overspreading us, and looks large enough on the IR satellite to be with us through sunrise.

We ended our night at 10UT (3am), and are going to get some sleep before our departure tomorrow around noon.

Our first official shutout...

-- RichardPogge - 19 Dec 2010
Topic revision: r10 - 22 Dec 2010, RichardGreen
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback